OpenStudy (afrodiddle):

which of the following is equal to \[\sqrt{\sqrt[3]{6}}\]

OpenStudy (afrodiddle):

6 ^1/6 6 ^1/3 6 ^2/3 6 ^3/2 Medal I will fan if you explain it to me :)

OpenStudy (afrodiddle):

@Austin6i6 Is this one 6 ^1/3 ?

OpenStudy (phi):

do you know how to write \[ \sqrt[3]{6} \] using exponents ?

OpenStudy (afrodiddle):

no, can you teach me?

OpenStudy (phi):

it's (unfortunately) memorization. the little 3 in front of the radical sign becomes the denominator (bottom number) in the fraction 1/3 that is the exponent.

OpenStudy (afrodiddle):

I thought so, but I was not quite sure.

OpenStudy (phi):

the "cube root" or \( \sqrt[3]{6} \) can be written as \( 6^{\frac{1}{3}} \)

OpenStudy (afrodiddle):

So I was right?

OpenStudy (phi):

that is part of the problem. so far you have \[ \sqrt{ \sqrt[3]{6} } = \sqrt{6^\frac{1}{3} }\]

OpenStudy (afrodiddle):

I think I am starting to understand a little more.

OpenStudy (phi):

a "normal" square root means the "little number" is understood to be 2 in other words \[ \sqrt{6^\frac{1}{3} }\] means the same thing as \[ \sqrt[2]{6^\frac{1}{3} } \]

OpenStudy (phi):

and the idea is that \[ \sqrt[2]{stuff} = (stuff)^\frac{1}{2}\]

OpenStudy (phi):

in other words, wrap up \( 6^\frac{1}{3} \) in parens (think, put it in a package) then put 1/2 as its exponent

OpenStudy (afrodiddle):

(6^1/3) ^ 1/2

OpenStudy (phi):

yes. then you remember this rule for exponents \[ \left( a^b\right)^c = a^{b\cdot c} \]

OpenStudy (phi):

in this case, multiply the two exponents to get a single new exponent.

OpenStudy (phi):

do you remember how to multiply fractions ?

OpenStudy (phi):

top * top divided by bottom * bottom

OpenStudy (afrodiddle):

1/6 ?

OpenStudy (phi):

yes. 1/6 is the new exponent so the answer is \[ 6^\frac{1}{6} \]

OpenStudy (afrodiddle):

I think I get it now, can I tag you in my later questions to confirm my answer?

OpenStudy (phi):

I'm signing off. but someone will check your work.

OpenStudy (afrodiddle):

Okay :) thanks for your help!

OpenStudy (phi):

yw